# Preprint Notice

### v2

@@ III-B1 The Packet Loss Paradigm
The nature of data transmission at high speeds differs fundamentally in both engineering and measurement from casual, end-user level data transfers. For high-speed networks, "packets" often vary in size. For example, for nodes connected to 10 Gbps and faster networks, it is common to employ "jumbo frame" Maximum Transmission Unit (MTU) of 9,014 or 9,018 bytes, rather than a constant 1,500-byte frame. Consequently, high-speed network device vendors typically specify performance using Bit Error Rates (BER) [28] rather than packet loss. While operators historically reported BERs more explicitly, modern practitioners typically must navigate complex monitoring dashboards such as the perfSONAR lookup service [29] to access such information.
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in [30]) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to -10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@
This result—a BER of $4\times 10^{-9}$, meaning roughly 4 erroneous bits in every billion—confirms that packet loss is negligible in well-engineered backbones.
This theoretical upper bound of network quality is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of [8] (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in [4]. Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This production-scale result provides practical evidence that packet loss did not constrain throughput.
@@ III-B2 The TCP CCA Paradigm
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" [31], which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”
It is a telling coincidence that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization [4]. This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work [4]. Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already addressed by RFC8312 [32]. Coincidentally, earlier work [4] described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 [33], titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
It is noteworthy that while Google researchers introduced the model-based BBR CCA in 2018—the same year CUBIC was already the default in widely used Linux operating system distributions such as CentOS 7—BBR has since progressed to version 3 [34]. However, the BBR implementation that ships with RHEL 9.6 remains version 1, which is notably more aggressive than later versions. Here, we compared the performance of BBR, CUBIC, and Reno (a classical TCP CCA algorithm that emerged in the 1990s) and provided the results in Figs 4-6. The following definitions and methodology are established before presenting the three figures.
@@
    • Streaming transfer: File size ranges from 4 MiB to 1 TiB; sizes are incremented in power of two, so there are 19 datasets
Each sweep iteration processed datasets sequentially from the smallest to the largest file size. This process was repeated for multiple iterations to gather statistical significance.
@@ III-B3 Practical Implications and Real-World Consequences
A June 12th article [37] titled “Chinese AI Companies Dodge U.S. Chip Curbs by Flying Suitcases of Hard Drives Abroad” detailed how four engineers transported 4.8 PB of training data by flying sixty hard drives from Beijing to Kuala Lumpur. This volume of data, sufficient for training large language models (LLMs) such as OpenAI’s ChatGPT, highlights the severe real-world constraints imposed by inefficient electronic data transfer.
A related article from Tom's Hardware [38] clarified that this physical transport was a “meticulously planned operation and took several months of preparation. In sharp contrast, the performance results illustrated in Figs. 4-6, combined with the entries in Table V, demonstrate that transferring this volume of data electronically, using a 100 Gbps link, would require at most one week.
@@ III-C Dedicated Private Lines Are Essential for High-Speed Testing
A common paradigm asserts that validating high-speed data transfer requires a dedicated, high-bandwidth WAN link. This belief creates a significant barrier, as such links are costly and complex to provision, requiring specialists to operate well. We confronted this directly in 2023 while developing data movement appliances with Intel Corp. Neither organization had a 100 Gbps WAN, raising a critical question: how could we rigorously verify performance without a production-grade link?
@@
A turning point was the discovery of a 2012 technical report, “Validating Linux Network Emulation” [41], which outlined a method using Linux's built-in traffic control (tc [42]) and network emulation (tc-netem [43]) tools. Recognizing the potential of this software-based approach to create a high-fidelity, 100 Gbps-capable testbed at a fraction of the cost, we implemented an enhanced version. The resulting testbed (Fig. 7) —built at the former Intel Swindon Lab in collaboration with Intel and HPE—consists of two HPE DL380 Gen 11 servers (configured as the appliances under test) connected via a switch to two Intel "latency servers." These servers use tc and tc-netem to impose precise delay, accurately simulating transcontinental links in an automated, remotely manageable setup.
@@ III-F Virtualization and cloud are universally useful
We recently received an inquiry about evaluating zx for a site in South Africa needing to transfer data to Europe at higher rates. The site's proposal to use a virtual machine (VM) immediately raised a red flag, illustrating a fundamental mismatch between virtualized environments and data movement performance.
@@
    Interrupt Overhead: Even with Single Root I/O Virtualization (SR-IOV) [48], interrupt handling in VMs incurs significant overhead due to interception and remapping by the hypervisor.
@@ III-F1 The Architectural Cost of Cloud Abstraction
While virtualization and cloud abstraction deliver significant benefits in manageability and elasticity, they impose substantial performance penalties on data movement. The canonical cloud data path—encumbered by layers of overheads such as hypervisor, virtualized storage, and HTTP/REST APIs—incurs a substantial performance inefficiency compared to a tuned bare-metal environment. This inefficiency is empirically observed in our tests to routinely around 30–50%. For instance, on 10 Gbps Elastic Compute Cloud (EC2) instances in 2019, we achieved a maximum of ~6 Gbps per VM, a 40% loss from line rate. More recently, as detailed in Fig. 11, the native cloud tooling (aws-cli) performed at a small fraction of the available capacity, while our co-designed appliance, even when partially firewalled, consistently attained a significantly higher percentage of the physical link's potential.
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads [49], fragment a simple file transfer into a multi-step coordination process—effectively a Rube Goldberg machine [50] for data movement.
@@ III-F2 A Critical Examination of Cloud Provider Metrics
  • Ignoring this latency [37] and [38] deserves scrutiny.
@@
Results are presented primarily in marketing form and omit key engineering details required for reproducibility (storage, network, system configuration).
@@ Results:
    Cloud’s Self-Imposed Bottleneck: Drastic underperformance of aws-cli highlights that the bottleneck is not the physical network, but the inefficient software stack and lack of end-to-end control.
@@ IV Methodology and Reproducibility
The zx data mover and associated deployment tooling represent proprietary commercial technology reserved for customers and partners. The provided materials allow independent verification of our benchmarking methodology without requiring access to our commercial software.
V Acknowledgment
@@
VI Experimental Setup
VI-A Validation Platforms
VI-B Data Availability and Licensing
In the spirit of reproducible research, all elements required to replicate the testing and benchmarking presented in this paper are provided in a public GitHub repository [22]. The zx software data mover is a commercially licensed product and is not available for public distribution.
VII Conclusions
The holistic co-design principle central to this work represents the culmination of a sustained engineering philosophy first articulated in our early architectural blueprints [58] and subsequently refined through practical implementation and public discourse [59], [60]. This maturing perspective underscores that consistent, high-performance data movement has always been fundamentally a systems engineering challenge rather than an isolated network optimization problem.
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup [3] and [4] shown in Fig. 12 (reproduced from [8] Appendix 6.1). This scale-out design used multiple older servers with aggregated 10Gbps interfaces, proving that the software architecture could distribute workload across nodes without cluster management overhead. Modern 2U2N servers, e.g., the SYS- [^2] BT-DNR [61] (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.

### v3

@@ III-B1 The Packet Loss Paradigm
The nature of data transmission at high speeds differs fundamentally in both engineering and measurement from casual, end-user level data transfers. For high-speed networks, "packets" often vary in size and are transmitted at rates where bit-level impairments dominate over per-packet events. For example, for nodes connected to 10 Gbps and faster networks, it is common to employ "jumbo frame" Maximum Transmission Unit (MTU) of 9,014 or 9,018 bytes, rather than a constant 1,500-byte frame. Consequently, high-speed network device vendors typically specify performance using Bit Error Rates (BER) [39] rather than packet loss. While operators historically reported BERs more explicitly, modern practitioners typically must extract such information indirectly via monitoring infrastructures such as the perfSONAR lookup service [43].
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in [58]) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@
This result—a BER of $4\times 10^{-9}$, meaning roughly 4 erroneous bits in every billion—confirms that packet loss is negligible in well-engineered backbones. Note that using jumbo frames would further reduce the implied BER.
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of [34] (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in [15]. Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.
@@ III-B2 The TCP CCA Paradigm
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" [6], which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization [15]. This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work [15]. Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 [48]. Coincidentally, earlier work [15] described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 [60], titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
It is noteworthy that while Google researchers introduced the model-based BBR CCA in 2018—the same year CUBIC was already the default in widely used Linux operating system distributions such as CentOS 7—BBR has since progressed to version 3 [26]. However, the BBR implementation that ships with RHEL 9.6 remains version 1, which is notably more aggressive than later versions. This distinction is relevant because BBRv1’s aggressiveness should, if anything, favor higher throughput in uncongested environments. Here, we compared the performance of BBR, CUBIC, and Reno (a classical TCP CCA algorithm that emerged in the 1990s) and provided the results in Figs 4-6. The following definitions and methodology are established before presenting the three figures.
@@
    • Streaming transfer sweep: File size ranges from 4 MiB to 1 TiB; sizes are incremented in power of two, so there are 19 datasets
Each sweep iteration processed datasets sequentially from the smallest to the largest file size. This process was repeated for multiple iterations to gather statistically stable results.
@@ III-B3 Practical Implications and Real-World Consequences
A June 12th article [59] titled “Chinese AI Companies Dodge U.S. Chip Curbs by Flying Suitcases of Hard Drives Abroad” detailed how four engineers transported 4.8 PB of training data by flying sixty hard drives from Beijing, China to Kuala Lumpur, Malaysia. This volume of data, sufficient for training large language models (LLMs) such as OpenAI’s ChatGPT, highlights the severe real-world constraints imposed by inefficient electronic data transfer.
A related article from Tom's Hardware [37] clarified that this physical transport was a “meticulously planned operation and took several months of preparation. In sharp contrast, the performance results illustrated in Figs. 4-6, combined with the entries in Table V, demonstrate that transferring this volume of data electronically, using a 100 Gbps link, would require at most one week of sustained transfer.
@@ III-C Dedicated Private Lines Are Essential for High-Speed Testing
A common paradigm asserts that validating high-speed data transfer requires a dedicated, high-bandwidth WAN link. This belief creates a significant barrier, as such links are costly and complex to provision, requiring specialists to operate well. We confronted this directly in 2023 while developing data movement appliances with Intel Corp. Neither organization had a 100 Gbps WAN, raising a critical question: how could we rigorously validate performance in the absence of a production-grade link?
@@
A turning point was the discovery of a 2012 technical report, “Validating Linux Network Emulation” [35], which outlined a method using Linux's built-in traffic control (tc [57]) and network emulation (tc-netem [56]) tools. Recognizing the potential of this software-based approach to create a high-fidelity, 100 Gbps capable testbed at a fraction of the cost, we implemented an enhanced version. This version extended functionality and automation beyond the original method. The resulting testbed (Fig. 7) —built at the former Intel Swindon Lab in collaboration with Intel and HPE—consists of two HPE DL380 Gen 11 servers (configured as the appliances under test) connected via a switch to two Intel "latency servers." These servers use tc and tc-netem to impose precise delay, accurately simulating transcontinental links in an automated, remotely manageable setup.
@@ III-F Virtualization and cloud are universally useful
We recently received an inquiry about evaluating zx for a site in South Africa needing to transfer data to Europe at higher rates. The site's proposal to use a virtual machine (VM) immediately raised concerns, illustrating a fundamental mismatch between virtualized environments and data movement performance.
@@
    Interrupt Overhead: Even with Single Root I/O Virtualization (SR-IOV) [41], interrupt handling in VMs incurs measurable overhead due to interception and remapping by the hypervisor.
@@ III-F1 The Architectural Cost of Cloud Abstraction
While virtualization and cloud abstraction deliver significant benefits in manageability and elasticity, they impose substantial performance penalties on data movement. The canonical cloud data path—encumbered by layers of overheads such as hypervisor, virtualized storage, and HTTP/REST APIs—incurs a substantial performance inefficiency compared to a tuned bare-metal environment. This inefficiency is empirically observed in our tests to routinely reach 30–50%. For instance, on 10 Gbps Elastic Compute Cloud (EC2) instances in 2019, we achieved a maximum of ~6 Gbps per VM, a 40% loss from line rate. More recently, as detailed in Fig. 11, the native cloud tooling (aws-cli) performed at a small fraction of the available capacity, while our co-designed appliance, even when partially firewalled, consistently attained a significantly higher percentage of the physical link's potential.
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads [3], fragment a simple file transfer into a multi-step coordination process—effectively a highly fragmented workflow [2] for data movement.
@@ III-F2 A Critical Examination of Cloud Provider Metrics
  • Treating this latency [59] and [37] deserves scrutiny.
@@
Results are presented primarily in marketing form and omit key engineering details required for reproducibility, including storage, network, system configuration.
@@ Results:
    Cloud’s Self-Imposed Bottleneck: The underperformance of aws-cli indicates that the bottleneck is not the physical network, but the inefficient software stack and limited end-to-end control.
@@ IV Methodology and Reproducibility
The zx data mover and associated deployment tooling represent proprietary commercial technology reserved for customers and partners. The provided materials allow independent verification of the benchmarking methodology and performance claims without requiring access to the commercial zx software.
V Deployments
V-A Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies [53, 14, 15, 28, 40, 17, 21].
V-B Data Availability and Licensing
In the spirit of reproducible research, all elements required to replicate the testing and benchmarking methodology presented in this paper are provided in a public GitHub repository [19]. The zx software data mover is a commercially licensed product and is not available for public distribution.
VI Conclusions
The holistic co-design principle central to this work represents the evolution of a decade-plus sustained and continuing engineering effort. This lineage traces back to foundational architectural blueprints developed at SLAC National Accelerator Laboratory [17] and subsequently validated through formal DOE/ESnet evaluation [34]. The architecture’s efficacy has been demonstrated through high-stakes production trials at the Linac Coherent Light Source (LCLS) [53] and through extensive public discourse [21], [18], [22]. This maturing perspective has recently been extended to modern hardware platforms, demonstrating that consistent, high-performance data movement can be seamlessly carried out on Data Processing Units (DPUs) [23], [62]. This maturing perspective underscores that consistent, high-performance data movement has always been fundamentally a systems engineering challenge rather than an isolated network optimization problem.
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup [14] and [15] shown in Fig. 12 (reproduced from [34] Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10Gbps mini-appliance or a 400Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR [51] (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.
The practical impact of this architectural approach extends beyond technical metrics: predictable, high-throughput data movement enables research, education, and industrial workflows at scales previously reserved for multi-million-dollar infrastructures, translating high-performance capability into tangible strategic and economic value.
VII Acknowledgment
@@
Chin Fang wishes to thank his colleague Riccardo Veraldi for valuable assistance with TCP congestion control and kernel optimization in 2024. While Chin was building the data movement appliances subsection III-C, Riccardo’s guidance contributed to the success of this work.